# A Simple First-Order Algorithm for Full-Rank Equality Constrained Optimization

### v1

@@ 3.2 Tangential steps
Our analysis hinges on the fact that first-order descent can be shown on the Lyapunov $\psi(x)$, both for tangential and normal steps, despite the fact that neither $\widehat{\lambda}(x_{k})$ or $\rho$ (which we still need to define) appears in the algorithm. We start by considering tangential steps.
@@
Suppose that AS.4–AS.7 hold. Then $\psi(x_{k_{\tau}+1})-\psi(x_{k_{\tau}})\leq-\alpha_{T,k_{\tau}}\|g_{T,k_{\tau}}\|^{2}+\eta\,\kappa_{\mbox{\tiny tan}}\,\alpha_{T,k_{\tau}}^{2}\|g_{T,k_{\tau}}\|^{2}.$ (3.2) where $\kappa_{\mbox{\tiny tan}}=\frac{1}{2}\Big(L_{L}+\rho L_{c}+\beta L_{\lambda}+L_{\lambda}L_{c}\Big).$
@@
    $\psi(x_{k_{\tau}+1})-\psi(x_{k_{\tau}})$ $=\underbrace{\psi_{\rho}(x_{k_{\tau}+1},\widehat{\lambda}(x_{k_{\tau}}))-\psi_{\rho}(x_{k_{\tau}},\widehat{\lambda}(x_{k_{\tau}}))}_{\Delta_{x}}$ (3.3)
    $\;\;\;\;\;\;\;\;\;\;\;\;+\underbrace{\psi_{\rho}(x_{k_{\tau}+1},\widehat{\lambda}(x_{k_{\tau}+1}))-\psi_{\rho}(x_{k_{\tau}+1},\widehat{\lambda}(x_{k_{\tau}}))}_{\Delta_{\lambda}}.$
@@
    $\Delta_{x}=-\nabla_{x}\psi_{\rho}(x_{k_{\tau}},\widehat{\lambda}(x_{k_{\tau}}))^{T}(\alpha_{T,k_{\tau}}g_{T,k_{\tau}})+r_{0}+\rho(\|c(x_{k_{\tau}+1})\|-\|c(x_{k_{\tau}})\|),$
@@
    $|r_{0}|\leq\frac{L_{L}}{2}\alpha_{T,k_{\tau}}^{2}\|g_{T,k_{\tau}}\|^{2}.$
@@
    $\|c(x_{k_{\tau}+1})\|=\|c(x_{k_{\tau}})-\alpha_{T,k_{\tau}}J_{k}g_{T,k_{\tau}}+r_{1}\|\leq\|c(x_{k_{\tau}})\|+\|r_{1}\|$
@@
    $\|r_{1}\|\leq\frac{L_{c}}{2}\alpha_{T,k_{\tau}}^{2}\|g_{T,k_{\tau}}\|^{2},$
@@
    $\|c(x_{k_{\tau}+1})\|-\|c(x_{k_{\tau}})\|\leq\frac{L_{c}}{2}\alpha_{T,k_{\tau}}^{2}\|g_{T,k_{\tau}}\|^{2}.$ (3.4)
    Successively using (2.8) and (2.9), we obtain that
    $\Delta_{x}$ $=-\alpha_{T,k_{\tau}}g_{T,k_{\tau}}^{T}g_{T,k_{\tau}}+r_{0}+\frac{\rho L_{c}}{2}\,\alpha_{T,k_{\tau}}^{2}\|g_{T,k_{\tau}}\|^{2}$ (3.5)
    $\leq-\alpha_{T,k_{\tau}}g_{T,k_{\tau}}^{T}g_{T,k_{\tau}}+\frac{1}{2}\left(L_{L}+\rho L_{c}\right)\,\alpha_{T,k_{\tau}}^{2}\|g_{T,k_{\tau}}\|^{2}$
    $\leq-\alpha_{T,k_{\tau}}\|g_{T,k_{\tau}}\|^{2}+\frac{1}{2}\left(L_{L}+\rho L_{c}\right)\,\alpha_{T,k_{\tau}}^{2}\|g_{T,k_{\tau}}\|^{2}.$
    Now, we may use the Lipschitz continuity of $\widehat{\lambda}$ and $c$ to deduce that
    $\Delta_{\lambda}$ $=c_{k_{\tau}+1}^{T}\big(\widehat{\lambda}_{k_{\tau}+1}-\widehat{\lambda}_{k_{\tau}}\big)$ (3.6)
    $=(c_{k_{\tau}+1}-c_{k_{\tau}})^{T}\big(\widehat{\lambda}_{k_{\tau}+1}-\widehat{\lambda}_{k_{\tau}}\big)+c_{k_{\tau}}^{T}\big(\widehat{\lambda}_{k_{\tau}+1}-\widehat{\lambda}_{k_{\tau}}\big)$
    $\leq\|c_{k_{\tau}}\|\,\|\widehat{\lambda}_{k_{\tau}+1}-\widehat{\lambda}_{k_{\tau}}\|+\|c_{k_{\tau}+1}-c_{k_{\tau}}\|\,\|\widehat{\lambda}_{k_{\tau}+1}-\widehat{\lambda}_{k_{\tau}}\|$
    $\leq L_{\lambda}\|c_{k_{\tau}}\|\,\alpha_{T,k_{\tau}}\|g_{T,k_{\tau}}\|+L_{\lambda}L_{c}\alpha_{T,k_{\tau}}^{2}\|g_{T,k_{\tau}}\|^{2},$
    Taking (2.12) into account gives that $\|c_{k_{\tau}}\|\leq\beta\alpha_{T,k_{\tau}}\|g_{T,k_{\tau}}\|$, we obtain that
    $\Delta_{\lambda}\leq\left(\beta L_{\lambda}+L_{\lambda}L_{c}\right)\alpha_{T,k_{\tau}}^{2}\|g_{T,k_{\tau}}\|^{2}.$
@@
    $\psi(x_{k_{\tau}+1})-\psi(x_{k_{\tau}})\leq$ $-\alpha_{T,k_{\tau}}\|g_{T,k_{\tau}}\|^{2}+\frac{1}{2}(L_{L}+\rho L_{c})\alpha_{T,k_{\tau}}^{2}\|g_{T,k_{\tau}}\|^{2}$
    $+\left(\beta L_{\lambda}+L_{\lambda}L_{c}\right)\alpha_{T,k_{\tau}}^{2}\|g_{T,k_{\tau}}\|^{2}$
@@
The bound (3.2) quantifies the effect of tangential steps on the Lyapunov function, and its right-hand side involves a first-order (descent) term and a second-order perturbation term. We now derive crucial bounds on these terms.
@@
Suppose that AS.2 and AS.5 hold. If we denote $\Gamma_{k_{\tau+1}}=\Gamma_{k_{\tau}}+\|g_{T,k_{\tau}}\|^{2}\;\;\mbox{ and }\;\;\alpha_{T,k_{\tau}}=\frac{\eta}{\sqrt{\Gamma_{k_{\tau+1}}+\varsigma}},$ then, for all $0\leq\tau_{0}\leq\tau_{1}$, $\sum_{\tau=\tau_{0}}^{\tau_{1}}\alpha_{T,k_{\tau}}\,\|g_{T,k_{\tau}}\|^{2}>\frac{\eta}{2\sqrt{2}}\,\sqrt{\Gamma_{k_{\tau_{1}+1}}}$ (3.7) $\sum_{\tau=\tau_{0}}^{\tau_{1}}\alpha_{T,k_{\tau}}^{2}\,\|g_{T,k_{\tau}}\|^{2}\leq\eta^{2}\,\log\left(\frac{\Gamma_{k_{\tau_{1}+1}}+\varsigma}{\Gamma_{k_{\tau_{0}}}+\varsigma}\right).$ (3.8)
  • Proof.    Let $w_{k_{\tau}+1}=\sqrt{\Gamma_{k_{\tau+1}}+\varsigma}$. The definition (2.11) implies that
    $\sum_{\tau=\tau_{0}}^{\tau_{1}}\alpha_{T,k_{\tau}}\,\|g_{T,k_{\tau}}\|^{2}$ $=\eta\sum_{\tau=\tau_{0}}^{\tau_{1}}\frac{\|g_{T,k_{\tau}}\|^{2}}{\sqrt{\Gamma_{k_{\tau+1}}+\varsigma}}$
    $>\eta\sum_{\tau=\tau_{0}}^{\tau_{1}}\frac{\|g_{T,k_{\tau}}\|^{2}}{w_{k_{\tau+1}}+w_{k_{\tau}}}$
    $=\eta\sum_{\tau=\tau_{0}}^{\tau_{1}}\frac{w_{k_{\tau+1}}^{2}-w_{k_{\tau}}^{2}}{w_{k_{\tau+1}}+w_{k_{\tau}}}$
    $=\eta\sum_{\tau=\tau_{0}}^{\tau_{1}}(w_{k_{\tau+1}}-w_{k_{\tau}}),$
    $=\eta(w_{k_{\tau_{1}+1}}-w_{k_{\tau_{0}}}),$
    Now observe that, using (3.1),
    $w_{k_{\tau_{1}+1}}-w_{k_{\tau_{0}}}$ $=\sqrt{\Gamma_{k_{\tau_{1}+1}}+\varsigma}-\sqrt{\Gamma_{k_{\tau_{0}}}+\varsigma}$
    $=\frac{\Gamma_{k_{\tau_{1}+1}}-\Gamma_{k_{\tau_{0}}}}{\sqrt{\Gamma_{k_{\tau_{1}+1}}+\varsigma}+\sqrt{\Gamma_{k_{\tau_{0}}}+\varsigma}}$
    $\geq\frac{\Gamma_{k_{\tau_{1}+1}}-\Gamma_{k_{\tau_{0}}}}{\sqrt{2\Gamma_{k_{\tau_{1}+1}}}+\sqrt{2\Gamma_{k_{\tau_{0}}}}}$
    $>\frac{\Gamma_{k_{\tau_{1}+1}}}{2\sqrt{2\Gamma_{k_{\tau_{1}+1}}}}=\frac{\sqrt{\Gamma_{k_{\tau_{1}+1}}}}{2\sqrt{2}}$
    which gives (3.7). Using the concavity and the increasing nature of the logarithm, we also have from (2.11) that
    $\sum_{\tau=\tau_{0}}^{\tau_{1}}\alpha_{T,k_{\tau}}^{2}\|g_{T,k_{\tau}}\|^{2}=\eta^{2}\!\sum_{\tau=\tau_{0}}^{\tau_{1}}\!\frac{\|g_{T,k_{\tau}}\|^{2}}{\Gamma_{k_{\tau+1}}+\varsigma}=\eta^{2}\!\sum_{\tau=\tau_{0}}^{\tau_{1}}\!\frac{\Gamma_{k_{\tau+1}}-\Gamma_{k_{\tau}}}{\Gamma_{k_{\tau+1}}+\varsigma}\leq\eta^{2}\!\sum_{\tau=\tau_{0}}^{\tau_{1}}\!\log(\Gamma_{k_{\tau+1}}+\varsigma)-\log(\Gamma_{k_{\tau}}+\varsigma),$
@@ 3.3 Normal steps
Having considered tangential steps, we now show that a decrease in the Lyapunov function $\psi$ may also be proved if a normal step is taken, provided the penalty parameter $\rho$ is chosen large enough. We start with a fairly simple observation.
@@
Suppose that AS.5 holds and that a normal step is used at iteration $k_{\nu}$. Then $\|c_{k_{\nu}+1}\|\leq\left(1-\kappa_{\mbox{\tiny nrm}}\sigma_{0}^{2}\right)\|c_{k_{\nu}}\|.$ (3.9)
@@
    $\|c_{k_{\nu}+1}\|-\|c_{k_{\nu}}\|\leq-\frac{\kappa_{\mbox{\tiny nrm}}\|J_{k_{\nu}}^{T}c_{k_{\nu}}\|^{2}}{\|c_{k_{\nu}}\|}\leq-\frac{\kappa_{\mbox{\tiny nrm}}\sigma_{0}^{2}\|c_{k_{\nu}}\|^{2}}{\|c_{k_{\nu}}\|}=-\kappa_{\mbox{\tiny nrm}}\sigma_{0}^{2}\|c_{k_{\nu}}\|.$
@@
Suppose that AS.3–AS.7 hold and that a normal step is used at iteration $k_{\nu}$. Define $\rho=\frac{1}{\kappa_{\rm nrm}\,\sigma_{0}^{2}}\Big[\;\kappa_{c}\,\theta\,\,L_{\lambda}\;+\;\frac{1}{2}\,\kappa_{c}\,\theta^{2}\,(L_{L}+L_{\lambda}L_{c})\;+\;\eta\;\Big]$ (3.10) Then we have that $\psi(x_{k_{\nu}+1})-\psi(x_{k_{\nu}})\leq-\eta\|c_{k_{\nu}}\|.$ (3.11)
@@
    $\leq(\nabla_{x}L(x_{k_{\nu}},\widehat{\lambda}_{k_{\nu}})^{T}s_{N,k_{\nu}}+r_{3}-\rho\kappa_{\mbox{\tiny nrm}}\sigma_{0}^{2}\|c_{k_{\nu}}\|$
    $\leq-\rho\kappa_{\mbox{\tiny nrm}}\sigma_{0}^{2}\|c_{k_{\nu}}\|+\frac{\theta^{2}L_{L}}{2}\|c_{k_{\nu}}\|^{2}.$
@@
    $\Delta_{\lambda}$ $=\psi_{\rho}(x_{k_{\nu}+1},\widehat{\lambda}_{k_{\nu}+1})-\psi_{\rho}(x_{k_{\nu}+1},\widehat{\lambda}_{k_{\nu}})$ (3.13)
@@
    $\leq L_{\lambda}\,\theta\|c_{k_{\nu}}\|^{2}+\frac{L_{\lambda}L_{c}\theta^{2}}{2}\|c_{k_{\nu}}\|^{2}$
@@
    $\psi(x_{k_{\nu}+1})-\psi(x_{k_{\nu}})\leq-\rho\kappa_{\mbox{\tiny nrm}}\sigma_{0}^{2}\|c_{k_{\nu}}\|+L_{\lambda}\,\theta\kappa_{c}\,\|c_{k_{\nu}}\|+\frac{1}{2}\left(L_{L}+L_{\lambda}L_{c}\right)\theta^{2}\kappa_{c}\|c_{k_{\nu}}\|.$
    The bound (3.11) then follows from (3.10).
    $\Box$

### v2

@@ 3.2 Tangential steps
Our analysis hinges on the fact that first-order descent can be shown on the Lyapounov $\psi(x)$, both for tangential and normal steps, despite the fact that neither $\widehat{\lambda}(x_{k})$ or $\rho$ (which we still need to define) appears in the algorithm. We start by considering tangential steps.
@@
Suppose that AS.4–AS.10 hold. Then $\mathbb{E}^{\tau}_{k}\!\left[\psi(x_{k+1})\right]-\psi(x_{k})\leq-\mathbb{E}^{\tau}_{k}\!\left[\alpha_{T,kx}\|g_{T,k}\|^{2}\right]+\eta\,\kappa_{\mbox{\tiny tan}}\,\mathbb{E}^{\tau}_{k}\!\left[\alpha_{T,k}^{2}\|g_{T,k}\|^{2}\right].$ (3.4) where $\kappa_{\mbox{\tiny tan}}=\frac{1}{2\eta}\Big(\kappa_{\mbox{\tiny dir}}+L_{L}+\rho L_{c}+\beta L_{\lambda}+L_{\lambda}L_{c}\Big).$
@@
    $\mathbb{E}^{\tau}_{k}\!\left[\psi(x_{k+1})\right]-\psi(x_{k})$ $=\underbrace{\mathbb{E}^{\tau}_{k}\!\left[\psi_{\rho}(x_{k+1},\widehat{\lambda}(x_{k}))\right]-\psi_{\rho}(x_{k},\widehat{\lambda}(x_{k}))}_{\Delta_{x}}$ (3.5)
    $\;\;\;\;\;\;\;\;\;\;\;\;+\underbrace{\mathbb{E}^{\tau}_{k}\!\left[\psi_{\rho}(x_{k+1},\widehat{\lambda}(x_{k+1}))-\psi_{\rho}(x_{k+1},\widehat{\lambda}(x_{k}))\right]}_{\Delta_{\lambda}}.$
@@
    $\Delta_{x}=-\mathbb{E}^{\tau}_{k}\!\left[(\nabla_{x}L_{\rho}(x_{k},\widehat{\lambda}(x_{k}))^{T}(\alpha_{T,k}g_{T,k})+r_{0}\right]+\rho(\mathbb{E}^{\tau}_{k}\!\left[\|c(x_{k+1})\|\right]-\|c(x_{k})\|),$
@@
    $\mathbb{E}^{\tau}_{k}\!\left[|r_{0}|\right]\leq\frac{L_{L}}{2}\mathbb{E}^{\tau}_{k}\!\left[\alpha_{T,k}^{2}\|g_{T,k}\|^{2}\right].$
@@
    $\mathbb{E}^{\tau}_{k}\!\left[\|c(x_{k+1})\|\right]=\mathbb{E}^{\tau}_{k}\!\left[\|c(x_{k})-\alpha_{T,k}J_{k}g_{T,k}+r_{1}\|\right]\leq\|c(x_{k})\|+\|r_{1}\|$
@@
    $\mathbb{E}^{\tau}_{k}\!\left[\|r_{1}\|\right]\leq\frac{L_{c}}{2}\mathbb{E}^{\tau}_{k}\!\left[\alpha_{T,k}^{2}\|g_{T,k}\|^{2}\right],$
@@
    $\mathbb{E}^{\tau}_{k}\!\left[\|c(x_{k+1})\|\right]-\|c(x_{k})\|\leq\frac{L_{c}}{2}\mathbb{E}^{\tau}_{k}\!\left[\alpha_{T,k}^{2}\|g_{T,k}\|^{2}\right].$ (3.6)
    Successively using (2.8), (2.9), Jensens’ inequality and the convexity of the absolute value, and AS.10, we obtain that
    $\Delta_{x}$ $=-\mathbb{E}^{\tau}_{k}\!\left[\alpha_{T,k}G_{T,k}^{T}g_{T,k}\right]+\mathbb{E}^{\tau}_{k}\!\left[r_{0}\right]+\frac{\rho L_{c}}{2}\,\mathbb{E}^{\tau}_{k}\!\left[\alpha_{T,k}^{2}\|g_{T,k}\|^{2}\right]$ (3.7)
    $\leq-\mathbb{E}^{\tau}_{k}\!\left[\alpha_{T,k}g_{T,k}^{T}g_{T,k}\right]+|\mathbb{E}^{\tau}_{k}\!\left[\alpha_{T,k}(G_{T,k}-g_{T,k})^{T}g_{T,k}\right]|+\frac{1}{2}\left(L_{l}+\rho L_{c}\right)\,\mathbb{E}^{\tau}_{k}\!\left[\alpha_{T,k}^{2}\|g_{T,k}\|^{2}\right]$
    $\leq-\mathbb{E}^{\tau}_{k}\!\left[\alpha_{T,k}g_{T,k}^{T}g_{T,k}\right]+\mathbb{E}^{\tau}_{k}\!\left[\alpha_{T,k}|G_{T,k}-g_{T,k})^{T}g_{T,k}|\right]+\frac{1}{2}\left(L_{l}+\rho L_{c}\right)\,\mathbb{E}^{\tau}_{k}\!\left[\alpha_{T,k}^{2}\|g_{T,k}\|^{2}\right]$
    $\leq-\mathbb{E}^{\tau}_{k}\!\left[\alpha_{T,k}\|g_{T,k}\|^{2}\right]+\frac{1}{2}\left(\kappa_{\mbox{\tiny dir}}+L_{L}+\rho L_{c}\right)\,\mathbb{E}^{\tau}_{k}\!\left[\alpha_{T,k}^{2}\|g_{T,k}\|^{2}\right].$
    Now, we may use the Lispchitz continuity of $\widehat{\lambda}$ and $c$ to deduce that
    $\Delta_{\lambda}$ $=\mathbb{E}^{\tau}_{k}\!\left[c_{k+1}^{T}\big(\widehat{\lambda}(x_{k+1})-\widehat{\lambda}(x_{k})\big)\right]$ (3.8)
    $=\mathbb{E}^{\tau}_{k}\!\left[(c_{k+1}-c_{k})^{T}\big(\widehat{\lambda}(x_{k+1})-\widehat{\lambda}(x_{k})\big)\right]+\mathbb{E}^{\tau}_{k}\!\left[c_{k}^{T}\big(\widehat{\lambda}(x_{k+1})-\widehat{\lambda}(x_{k})\big)\right]$
    $\leq\mathbb{E}^{\tau}_{k}\!\left[\|c_{k}\|\,\|\widehat{\lambda}(x_{k+1})-\widehat{\lambda}(x_{k})\|\right]+\mathbb{E}^{\tau}_{k}\!\left[\|c_{k+1}-c_{k}\|\,\|\widehat{\lambda}(x_{k+1})-\widehat{\lambda}(x_{k})\|\right]$
    $\leq L_{\lambda}\mathbb{E}^{\tau}_{k}\!\left[\|c_{k}\|\alpha_{T,k}\|g_{T,k}\|\right]+L_{\lambda}L_{c}\mathbb{E}^{\tau}_{k}\!\left[\alpha_{T,k}^{2}\|g_{T,k}\|^{2}\right],$
    Taking into account the fact that, for $k\in\{k_{\tau}\}$, (2.12) gives that $\|c_{k}\|\leq\beta\alpha_{T,k}\|g_{T,k}\|$, we obtain that
    $\Delta_{\lambda}\leq\left(\beta L_{\lambda}+L_{\lambda}L_{c}\right)\mathbb{E}^{\tau}_{k}\!\left[\alpha_{T,k}^{2}\|g_{T,k}\|^{2}\right].$
@@
    $\mathbb{E}^{\tau}_{k}\!\left[\psi(x_{k_{\tau}+1})\right]-\psi(x_{k_{\tau}})\leq$ $-\mathbb{E}^{\tau}_{k}\!\left[\alpha_{T,k_{\tau}}\|g_{T,k_{\tau}}\|^{2}\right]+\frac{1}{2}(\kappa_{\mbox{\tiny dir}}+L_{L}+\rho L_{c})\mathbb{E}_{k}\!\left[\alpha_{T,k_{\tau}}^{2}\|g_{T,k_{\tau}}\|^{2}\right]$
    $+\left(\beta L_{\lambda}+L_{\lambda}L_{c}\right)\mathbb{E}^{\tau}_{k}\!\left[\alpha_{T,k}^{2}\|g_{T,k}\|^{2}\right]$
@@
The bound (3.4) quantifies the effect of tangential steps on the Lyapounov function, and its right-hand side involves a first-order (descent) term and a second-order perturbation term. We now derive crucial bounds on these terms.
@@
Suppose that AS.2 and AS.5 hold. If we denote $\Gamma_{k_{\tau+1}}=\Gamma_{k_{\tau}}+\|g_{T,k_{\tau}}\|^{2},\quad\alpha_{T,k_{\tau}}=\frac{\eta}{\sqrt{\Gamma_{k_{\tau}}+\varsigma}},\quad w_{k_{\tau}}=\sqrt{\Gamma_{k_{\tau}}+\varsigma^{2}},\quad$ then, for all $0\leq\tau_{0}\leq\tau_{1}$, $\sum_{\tau=\tau_{0}}^{\tau_{1}}\alpha_{T,k_{\tau}}\,\|g_{T,k_{\tau}}\|^{2}>\eta\big(w_{k_{\tau_{1}}}-w_{k_{\tau_{0}-1}}\big),\$(3.9) $\sum_{\tau=\tau_{0}}^{\tau_{1}}\alpha_{T,k_{\tau}}^{2}\,\|g_{T,k_{\tau}}\|^{2}\leq\eta^{2}\,\log\left(\frac{\Gamma_{k_{\tau_{1}}}+\varsigma}{\Gamma_{k_{\tau_{0}-1}}+\varsigma}\right),$ (3.10)
  • Proof.    The definition (2.11) implies that, for every realization,
    $\sum_{\tau=\tau_{0}}^{\tau_{1}}\alpha_{T,k_{\tau}}\,\|g_{T,k_{\tau}}\|^{2}$ $=\eta\sum_{\tau=\tau_{0}}^{\tau_{1}}\frac{\|g_{T,k_{\tau}}\|^{2}}{\sqrt{\Gamma_{k_{\tau}}+\varsigma}}$
    $>\eta\sum_{\tau=\tau_{0}}^{\tau_{1}}\frac{\|g_{T,k_{\tau}}\|^{2}}{w_{k_{\tau}}+w_{k_{\tau}-1}}$
    $=\eta\sum_{\tau=\tau_{0}}^{\tau_{1}}\frac{w_{k_{\tau}}^{2}-w_{k_{\tau-1}}^{2}}{w_{k_{\tau}}+w_{k_{\tau-1}}}$
    $=\eta\sum_{\tau=\tau_{0}}^{\tau_{1}}(w_{k_{\tau}}-w_{k_{\tau-1}}),$
    which gives (3.9) after taking the total expectation of both sides of the inequality. Using the concavity and the increasing nature of the logarithm, we also have from (2.11) that
    $\sum_{\tau=\tau_{0}}^{\tau_{1}}\alpha_{T,k_{\tau}}^{2}\|g_{T,k_{\tau}}\|^{2}=\eta^{2}\!\sum_{\tau=\tau_{0}}^{\tau_{1}}\!\frac{\|g_{T,k_{\tau}}\|^{2}}{\Gamma_{k_{\tau}}+\varsigma}=\eta^{2}\!\sum_{\tau=\tau_{0}}^{\tau_{1}}\!\frac{\Gamma_{k_{\tau}}-\Gamma_{k_{\tau-1}}}{\Gamma_{k_{\tau}}+\varsigma}\leq\eta^{2}\!\sum_{\tau=\tau_{0}}^{\tau_{1}}\!\log(\Gamma_{k_{\tau}}+\varsigma)-\log(\Gamma_{k_{\tau-1}}+\varsigma),$
@@ 3.3 Normal steps
Having considered tangential steps, we now show that a decrease in the Lyapounov function $\psi$ may also be proved if a normal step is taken, provided the penalty parameter $\rho$ is chosen large enough. Note that the normal step is deterministic. We start with a fairly simple observation.
@@
Suppose that AS.5 holds and that a normal step is used at iteration $k_{\nu}$. Then $\|c_{k_{\nu}+1}\|\leq\left(1-\kappa_{\mbox{\tiny nrm}}\kappa_{J}^{2}\right)\|c_{k_{\nu}}\|.$ (3.11)
@@
    $\|c_{k_{\nu}+1}\|-\|c_{k_{\nu}}\|\leq-\frac{\kappa_{\mbox{\tiny nrm}}\|J_{k_{\nu}}^{T}c_{k_{\nu}}\|^{2}}{\|c_{k_{\nu}}\|}\leq-\frac{\kappa_{\mbox{\tiny nrm}}\kappa_{J}^{2}\|c_{k_{\nu}}\|^{2}}{\|c_{k_{\nu}}\|}=-\kappa_{\mbox{\tiny nrm}}\kappa_{J}^{2}\|c_{k_{\nu}}\|.$
@@
Suppose that AS.4–AS.7 hold and that a normal step is used at iteration $k_{\nu}$. Define $\rho=\frac{\kappa_{c}\theta}{\kappa_{\mbox{\tiny nrm}}\kappa_{J}}\left[L_{\lambda}+\frac{\theta}{2}(L_{L}+L_{\lambda}L_{c})+\eta\right].$ (3.12) Then we have that $\psi(x_{k_{\nu}+1})-\psi(x_{k_{\nu}})\leq-\eta\|c_{k_{\nu}}\|.$ (3.13)
@@
    $\leq(\nabla_{x}L(x_{k_{\nu}},\widehat{\lambda}_{k_{\nu}})^{T}s_{N,k_{\nu}}+r_{3}-\rho\kappa_{\mbox{\tiny nrm}}\kappa_{J}^{2}\|c_{k_{\nu}}\|$
    $\leq-\rho\kappa_{\mbox{\tiny nrm}}\kappa_{J}^{2}\|c_{k_{\nu}}\|+\frac{\theta^{2}L_{L}}{2}\|J_{k_{\nu}}^{T}c_{k_{\nu}}\|^{2}.$
@@
    $\Delta_{\lambda}$ $=\psi_{\rho}(x_{k_{\nu}+1},\widehat{\lambda}(x_{k_{\nu}+1}))-\psi_{\rho}(x_{k_{\nu}+1},\widehat{\lambda}(x_{k_{\nu}}))$ (3.15)
@@
    $\leq L_{\lambda}\,\theta\|J_{k_{\nu}}^{T}c_{k_{\nu}}\|\,\|c_{k_{\nu}}\|+\frac{L_{\lambda}L_{c}\theta^{2}}{2}\|J_{k_{\nu}}^{T}c_{k_{\nu}}\|^{2}$
@@
    $\psi(x_{k_{\nu}+1})-\psi(x_{k_{\nu}})$ $\leq-\rho\kappa_{\mbox{\tiny nrm}}\kappa_{J}^{2}\|c_{k_{\nu}}\|+L_{\lambda}\,\theta\|J_{k_{\nu}}^{T}c_{k_{\nu}}\|\,\|c_{k_{\nu}}\|+\left(\frac{\theta^{2}L_{L}}{2}+\frac{L_{\lambda}L_{c}\theta^{2}}{2}\right)\|J_{k_{\nu}}^{T}c_{k_{\nu}}\|^{2}$
    $\leq-\rho\kappa_{\mbox{\tiny nrm}}\kappa_{J}^{2}\|c_{k_{\nu}}\|+L_{\lambda}\,\theta\kappa_{J}\kappa_{c}\,\|c_{k_{\nu}}\|+\left(\frac{\theta^{2}L_{L}}{2}+\frac{L_{\lambda}L_{c}\theta^{2}}{2}\right)\kappa_{J}^{2}\kappa_{c}\|c_{k_{\nu}}\|.$
    The bound (3.13) then follows from (3.12). $\Box$